# Non-Markovian Reduced Models to Unravel Transitions in Non-equilibrium Systems

### v1

@@ III.2 Non-Markovian path-dependent coefficients: Efficient simulation
Our stochastic parameterization given by (45) contains random coefficients $Z^{n}_{\tau}({\bm{\lambda}},t;\omega)$, each of them involving an integral of the history of the Brownian motion making them non-Markovian in the sense of [108], i.e. depending on the noise path history. We present below an efficient mean to compute these random coefficients by solving auxiliary ODEs with path-dependent coefficients, avoiding this way the cumbersome computation of integrals over noise paths that would need to be updated at each time $t$. This approach is used for the following purpose:
@@
According to (41), the computation of $Z^{n}_{\tau}$ boils down to the computation of the integral
@@
where $\tau>0$, and $j\geq m_{c}+1$ due to assumption (H). We only need to consider the case that $\kappa<0$, since we consider here only cases in which the unstable modes are included within the reduced state space $H_{\mathfrak{c}}$.
@@
Now, to determine $Z^{n}_{\tau}({\bm{\lambda}},t;\omega)$ in (41) it is sufficient to observe that
$Z^{n}_{\tau}({\bm{\lambda}},t;\omega)=\sigma_{n}\lambda_{n}I_{\tau}^{n}(t,\omega;\lambda_{n}).$ (49)
The random coefficient $Z^{n}_{\tau}({\bm{\lambda}},t;\omega)$ is thus computed by using (49) after $I_{\tau}^{n}(t,\omega;\lambda_{n})$ is computed by solving, forward in time, the corresponding RDE of the form (47) with initial datum
@@
This way, we compute only the integral (50) and then propagate it through Eq. (47) to evaluate $Z^{n}_{\tau}({\bm{\lambda}},t;\omega)$, instead of computing for each $t$ the integral in the definition (41). The resulting procedure is thus much more efficient to handle numerically than a direct computation of $Z^{n}_{\tau}({\bm{\lambda}},t;\omega)$ based on the integral formulation (41) which would involve a careful and computationally more expensive quadrature at each time-step. Note that a similar treatment has been adopted in [30, Chap. 5.3] for the computation of the time-dependent coefficients arising in stochastic parameterizations of SPDEs driven by linear multiplicative noise, and in [28, Sec. III.C.2] for the case of non-autonomous forcing.
@@ III.3 Data-informed optimization and training path
Sections V and VI below illustrate how this single-path training strategy can be used efficiently to predict from the corresponding reduced systems the statistical behavior as time and/or the noise’s path is varied. The next section delves into the justification of this variational approach, and the theoretical characterization of the notion of stochastic OPM by relying on ergodic arguments. The reader interested by applications can jump to Sections V and VI.
@@
IV Non-Markovian Optimal Parametrizations and Invariant Measures
@@ IV.1 Theoretical insights
A random probability measure $\rho_{\omega}$ that satisfies (53) is called below a (random) statistical equilibrium. Recall that $\theta_{t}$ here denotes the standard ergodic transformation on the set of Wiener path $\Omega$ defined through the helix identity [34, Def. 2.3.6], $W_{s}(\theta_{t}\omega)=W_{t+s}(\omega)-W_{t}(\omega)$, $\omega$ in $\Omega$.
@@
see [114, Prop. 4.2]; see also [34, Theorems 1.7.2 and 2.3.45].
@@
is a bounded martingale, and therefore converges $\mathbb{P}$-a.s by the Doob’s first martingale convergence theorem; see e.g. [117, Thm. C.5 p. 302]. This implies, in particular, the $\mathbb{P}$-almost sure convergence of the measure-valued random variable $\omega\mapsto S(t,\theta_{-t}\omega)\mu$ in the topology of weak convergence of $Pr_{\Omega}(H)$, which in turn implies the convergence in the narrow topology of $Pr_{\Omega}(H)$; see [118, Chap. 3].
@@
is an invariant measure of the Markov semigroup $P_{t}$; see [119, Thm. 1.10.1].
@@
In case a random attractor $\mathcal{A}(\omega)$ exists, since the latter is a forward invariant compact random set, the Markovian random measure $\rho_{\omega}$ must be supported by $\mathcal{A}(\omega)$; see [118, Thm. 6.17] (see also [120, Prop. 4.5]). We conclude thus —  in the case of a unique ergodic measure for $P_{t}$ —  to the existence of a unique Markov $S$-invariant random measure $\rho_{\omega}$ supported by $\mathcal{A}$, and that all other $S$-invariant random measures (also necessarily supported by $\mathcal{A}$) are non-Markovian.
RDSs where the future and past $\sigma$-algebras are independent, are generated for many stochastic systems in practice. This is the case for instance of a broad class of stochastic differential equations (SDEs); see [34, Sect. 2.3]. The problem of generation of white noise RDSs in infinite dimension is much less clarified and it is beyond the scope of this article to address this question.
@@
For a proof of Lemma IV.1 see that of [115, Lemma 2.5].
@@
Given a parameterization $\Phi:\Omega\times H_{\mathfrak{c}}\rightarrow H_{\mathfrak{s}}$ that is measurable (in the measure-theoretic sense), and a mild solution $u$ of Eq. (36), the pullback parameterization defect associated with $\Phi$ over the interval $[0,T]$ and for a noise-path $\omega$, is defined as:
$\mathcal{Q}_{T}(\omega,\Phi)=\frac{1}{T}\int_{0}^{T}\left\lVert u_{\mathfrak{s}}(t,\theta_{-t}\omega)-\Phi(\omega,u_{\mathfrak{c}}(t,\theta_{-t}\omega))\right\rVert^{2}\,\mathrm{d}t.$ (60)
@@
The random measure $\mathfrak{m}_{\omega}$ allows us to consider the following functional space $\mathcal{X}=L^{2}_{\mathfrak{m}}(\Omega\times H_{\mathfrak{c}};H_{\mathfrak{s}})$ of parameterizations, defined as
$\mathcal{X}=\bigg{\{}\Phi:\Omega\times H_{\mathfrak{c}}\rightarrow H_{\mathfrak{s}}\,\textrm{measurable}\;\Big{|}$ (62)
@@
namely the Hilbert space constituted by $H_{\mathfrak{s}}$-valued random functions of the resolved variables $X$ in $H_{\mathfrak{c}}$, that are square-integrable with respect to $\mathfrak{m}_{\omega}$, almost surely.
@@
Let us denote by $\rho_{\omega}$ the weak-limit of $\mu$ defined in (52) for case (i), and by $\rho_{\omega}$ the statistical equilibrium ensured by Lemma IV.1, in case (ii). We denote by $\rho_{\omega}^{X}$ the disintegration of $\rho_{\omega}$ on the small-scale subspace $H_{\mathfrak{s}}$, conditioned on the coarse-scale variable $X$.
@@
$\underset{\Phi\in\mathcal{X}}{\min}\int_{\Omega}\int_{(X,Y)\in H_{\mathfrak{c}}\times H_{\mathfrak{s}}}\left\lVert Y-\Phi(\omega,X)\right\rVert^{2}\,\mathrm{d}\rho_{\omega}(X,Y)\,\mathrm{d}\mathbb{P}(\omega),$ (64)
@@
$\Phi^{\ast}(\omega,X)=\int_{H_{\mathfrak{s}}}Y\,\mathrm{d}\rho_{\omega}^{X}(Y),\;\;X\in H_{\mathfrak{c}},\;\;\;\mathbb{P}\mbox{-}\rm{a.s.}$ (65)
@@
    The probability measure $\mu$ therein by the probability measure $\rho$ on $\Omega\times H$ that is naturally associated with the family of random measures $\rho_{\omega}$ and whose marginal on $\Omega$ is given by $\mathbb{P}$; see [114, Prop. 3.6],
@@
    The function space $L^{2}_{\mu}(H_{\mathfrak{c}}\times H_{\mathfrak{s}};H_{\mathfrak{s}})$ therein by the function space $\mathcal{E}=L^{2}_{\rho}(\Omega\times H_{\mathfrak{c}}\times H_{\mathfrak{s}};H_{\mathfrak{s}})$.
By applying to the ambient Hilbert space $\mathcal{E}$, the standard projection theorem onto closed convex sets [121, Theorem 5.2], one defines (given $\Pi_{\mathfrak{c}}$) the conditional expectation $\mathbb{E}_{\rho}[g|\Pi_{\mathfrak{c}}]$ of $g$ as the unique function in $\mathcal{E}$ that satisfies the inequality
@@
$\mathbb{E}_{\rho}[g|\Pi_{\mathfrak{c}}](\omega,X)=\int_{H_{\mathfrak{s}}}g(\omega,X,Y)\,\mathrm{d}\rho_{\omega,X}(Y),$ (69)
@@ IV.2 Non-Markovian optimal reduced model, and conditional expectation
$\mathbb{E}_{\rho_{\theta_{t}\omega}}[G|\Pi_{\mathfrak{c}}](X)=\int_{Y\in H_{\mathfrak{s}}}G(X+Y)\,\mathrm{d}\rho_{\theta_{t}\omega}^{X}(Y),\;X\in H_{\mathfrak{c}},$
@@
$\langle G_{2}(Y,Y)\rangle_{\rho_{\theta_{t}\omega}^{X}}=\int_{Y\in H_{\mathfrak{s}}}\Pi_{\mathfrak{c}}G_{2}(Y,Y)\,\mathrm{d}\rho_{\theta_{t}\omega}^{X}(Y),$ (72)
@@
Sometimes, we will simply refer to Eq. (73) as the (random) conditional expectation.
@@ IV.3 Practical implications
Our numerical results demonstrate that this class $\mathcal{P}$ effectively models and approximates the true non-Markovian terms in the optimal reduced model (Eq. (73)) across various scenarios. After learning an optimal parameterization within class $\mathcal{P}$ using a single noise realization, the resulting approximation of the optimal reduced model typically becomes an ODE system with coefficients dependent on the specific noise path. These path-dependent coefficients capture the interactions between noise and nonlinearities, as exemplified by the reduced models Eqns. (83) and (LABEL:Eq_reduced_S_shaped) below, and their ability to predict noise-induced transitions within the reduced state spaces.
@@ V.1 Noise-induced transitions in a stochastic Allen-Cahn model
with the potential $V$ given by the standard double-well function $V(u)=(u^{2}-1)^{2}/4$. The sACE provides a general framework for studying pattern formation and interface dynamics in systems ranging from liquid crystals and ferromagnets to tumor growth and cell membranes [1, 46].
Its universality makes it relevant across diverse fields, contributing to a unified understanding of these complex phenomena. By incorporating the spatio-temporal noise term, $\,\mathrm{d}{\bm{W}_{t}}$, into the deterministic Allen-Cahn equation, the sACE captures the inherent fluctuation and uncertainty present in real systems [45]. This allows for a more realistic description of transitions between phases, providing valuable insights beyond what pure deterministic models can offer. Near critical points, where phases coexist, the sACE reveals the delicate balance between deterministic driving forces and stochastic fluctuations. Studying these transitions helps us understand critical phenomena in diverse systems, from superfluid transitions to magnetization reversal in magnets [1, 45].
@@
As mentioned in Introduction, the derivation of reduced equations able to capture stochastic transitions in sACEs has been principally studied near the onset of instability [48, 49, 50, 51, 52].
In this study, we consider the sACE placed after several bifurcations have taken place where multiple steady states coexist; see Fig. 2. It corresponds to the parameter regime
@@
More specifically, the domain size $L$ is chosen so that when $\sigma=0$ the trivial steady state has experienced three successive bifurcations leading thus to a total of seven steady states with two stable nodes ($\phi_{1}^{+}$ and $\phi_{1}^{-}$) and five saddle points (00, $\phi_{2}^{a}$, $\phi_{2}^{b}$, $\phi_{3}^{a}$, and $\phi_{3}^{b}$). These steady states are characterized here by their zeros: the saddle states are of sign-change type while the stable ones are of constant sign; see [124, 129]. Note also that the same conclusion holds for the case with either Neumann boundary conditions or periodic boundary conditions; see e.g. [127, Figure 2]. In our notation, $a$ and $b$ denote the string made of sign changes over $(0,L)$, with leftmost symbol in $a$ to be $+$. For instance, for $\phi_{3}^{a}$ (resp. $\phi_{3}^{b}$), $a=\{+,-,+\}$ (resp. $b=\{-,+,-\}$), while $a=\{+,-\}$ (resp. $b=\{-,+\}$) for $\phi_{2}^{a}$ (resp. $\phi_{2}^{b}$).
@@
Figure 3: Typical and rare transition paths for Eq. (74). One path connecting $u_{0}=0$ to a sign-changing metastable state (typical-left panel), and one path connecting $u_{0}=0$ to a constant-sign metastable state (rare event-right panel). Are shown also the large ($u_{\mathfrak{c}}$) and small ($u_{\mathfrak{s}}$) scale contributions.
Noteworthy is the visual rendering of the sojourn time of the solution path shown in Fig. 2B. It is more substantial in the neighborhood of $\phi_{2}^{a}$ than in that of $\phi_{1}^{+}$. It is a depiction of a solution path to Eq. (74) over $[0,T]$: most of the trajectory spends time near a sign-change metastable state ($\phi_{2}^{a}$) rather than to the constant-sign one ($\phi_{1}^{+}$). Large ensemble simulations of the sACE for the parameter regime (76) reveals that what is shown for a single path here is actually observed at the final time $t=T$, when the noise path is varied: the sign-change states $\phi_{2}^{a/b}$ are the rule whereas the constant-sign states ϕ1±superscriptsubscriptitalic-$\phi_{1}^{\pm}$ are the exceptions, corresponding here to rare events. Figure 3 shows two solution paths in the space-time domain connecting $u_{0}=0$ to either a sign-changing metastable state (typical) or a constant-sign metastable state (rare event). In Section V.3 below we shall discuss the mechanisms behind these patterns and their transition phenomenology, and the extent to which non-Markovian reduced models described below are able to capture it.

### v2

@@ III.2 Non-Markovian path-dependent coefficients: Efficient simulation
Our stochastic parameterization given by (47) contains random coefficients $Z^{n}_{\tau}(t;\omega)$, each of them involving an integral of the history of the Brownian motion making them non-Markovian in the sense of [110], i.e. depending on the noise path history. We present below an efficient mean to compute these random coefficients by solving auxiliary ODEs with path-dependent coefficients, avoiding this way the cumbersome computation of integrals over noise paths that would need to be updated at each time $t$. This approach is used for the following purpose:
@@
According to (41), the computation of $Z^{n}_{\tau}$ boils down to the computation of the random integral
@@
where $\tau>0$, and $j\geq m_{c}+1$ due to assumption (H). We only need to consider the case that $\kappa<0$, since we consider here only cases in which the unstable modes are included within the space $H_{\mathfrak{c}}$ of resolved scales.
@@
Now, to determine $Z^{n}_{\tau}(t;\omega)$ in (41) it is sufficient to observe that
$Z^{n}_{\tau}(t;\omega)=\sigma_{n}\lambda_{n}I_{\tau}^{n}(t,\omega;\lambda_{n}).$ (51)
The random coefficient $Z^{n}_{\tau}(t;\omega)$ is thus computed by using (51) after $I_{\tau}^{n}(t,\omega;\lambda_{n})$ is computed by solving, forward in time, the corresponding RDE of the form (49) with initial datum
@@
This way, we compute only the integral (52) and then propagate it through Eq. (49) to evaluate $Z^{n}_{\tau}(t;\omega)$, instead of computing for each $t$ the integral in the definition (41). The resulting procedure is thus much more efficient to handle numerically than a direct computation of $Z^{n}_{\tau}(t;\omega)$ based on the integral formulation (41) which would involve a careful and computationally more expensive quadrature at each time-step. Note that a similar treatment has been adopted in [29, Chap. 5.3] for the computation of the time-dependent coefficients arising in stochastic parameterizations of SPDEs driven by linear multiplicative noise, and in [28, Sec. III.C.2] for the case of non-autonomous forcing.
@@ III.3 Data-informed optimization and training path
Sections V and VI below illustrate how this single-path training strategy can be used efficiently to predict from the corresponding reduced systems the statistical behavior as time and/or the noise’s path is varied. The next section delves into the justification of this variational approach, and the theoretical characterization of the notion of stochastic OPM by relying on ergodic arguments. The reader interested in applications can jump to Sections V and VI.
@@
IV Non-Markovian Optimal Parameterizations and Invariant Measures
@@ IV.1 Theoretical insights
A random probability measure $\rho_{\omega}$ that satisfies (55) is called below a (random) statistical equilibrium. Recall that $\theta_{t}$ here denotes the standard ergodic transformation on the set of Wiener path $\Omega$ defined through the helix identity [50, Def. 2.3.6], $W_{s}(\theta_{t}\omega)=W_{t+s}(\omega)-W_{t}(\omega)$, $\omega$ in $\Omega$.
@@
see [117, Prop. 4.2]; see also [50, Theorems 1.7.2 and 2.3.45].
@@
is a bounded martingale, and therefore converges $\mathbb{P}$-a.s by the Doob’s first martingale convergence theorem; see e.g. [111, Thm. C.5 p. 302]. This implies, in particular, the $\mathbb{P}$-almost sure convergence of the measure-valued random variable $\omega\mapsto S(t,\theta_{-t}\omega)\mu$ in the topology of weak convergence of $Pr_{\Omega}(H)$, which in turn implies the convergence in the narrow topology of $Pr_{\Omega}(H)$; see [120, Chap. 3].
@@
is an invariant measure of the Markov semigroup $P_{t}$; see [121, Thm. 1.10.1].
@@
In case a random attractor $\mathcal{A}(\omega)$ exists, since the latter is a forward invariant compact random set, the Markovian random measure $\rho_{\omega}$ must be supported by $\mathcal{A}(\omega)$; see [120, Thm. 6.17] (see also [122, Prop. 4.5]). We conclude thus —  in the case of a unique ergodic measure for $P_{t}$ —  to the existence of a unique Markov $S$-invariant random measure $\rho_{\omega}$ supported by $\mathcal{A}$, and that all other $S$-invariant random measures (also necessarily supported by $\mathcal{A}$) are non-Markovian.
RDSs where the future and past $\sigma$-algebras are independent, are generated for many stochastic systems in practice. This is the case for instance of a broad class of stochastic differential equations (SDEs); see [50, Sect. 2.3]. The problem of generation of white noise RDSs in infinite dimension is much less clarified and it is beyond the scope of this article to address this question.
@@
For a proof of Lemma IV.1 see that of [118, Lemma 2.5].
@@
Given a parameterization $\Phi:\Omega\times H_{\mathfrak{c}}\rightarrow H_{\mathfrak{f}}$ that is measurable (in the measure-theoretic sense), and a mild solution $u$ of Eq. (36), the pullback parameterization defect associated with $\Phi$ over the interval $[0,T]$ and for a noise-path $\omega$, is defined as:
$\mathcal{Q}_{T}(\omega,\Phi)=\frac{1}{T}\int_{0}^{T}\left\lVert u_{\mathfrak{f}}(t,\theta_{-t}\omega)-\Phi(\omega,u_{\mathfrak{c}}(t,\theta_{-t}\omega))\right\rVert^{2}\,\mathrm{d}t.$ (62)
@@
The random measure $\mathfrak{m}_{\omega}$ allows us to consider the following functional space $\mathcal{X}=L^{2}_{\mathfrak{m}}(\Omega\times H_{\mathfrak{c}};H_{\mathfrak{f}})$ of parameterizations, defined as
$\mathcal{X}=\bigg{\{}\Phi:\Omega\times H_{\mathfrak{c}}\rightarrow H_{\mathfrak{f}}\,\textrm{measurable}\;\Big{|}$ (64)
@@
namely the Hilbert space constituted by $H_{\mathfrak{f}}$-valued random functions of the resolved variables $X$ in $H_{\mathfrak{c}}$, that are square-integrable with respect to $\mathfrak{m}_{\omega}$, almost surely.
@@
Let us denote by $\rho_{\omega}$ the weak-limit of $\mu$ defined in (54) for case (i), and by $\rho_{\omega}$ the statistical equilibrium ensured by Lemma IV.1, in case (ii). We denote by $\rho_{\omega}^{X}$ the disintegration of $\rho_{\omega}$ on the small-scale subspace $H_{\mathfrak{f}}$, conditioned on the coarse-scale variable $X$.
@@
$\underset{\Phi\in\mathcal{X}}{\min}\int_{\Omega}\int_{(X,Y)\in H_{\mathfrak{c}}\times H_{\mathfrak{f}}}\left\lVert Y-\Phi(\omega,X)\right\rVert^{2}\,\mathrm{d}\rho_{\omega}(X,Y)\,\mathrm{d}\mathbb{P}(\omega),$ (66)
@@
$\Phi^{\ast}(\omega,X)=\int_{H_{\mathfrak{f}}}Y\,\mathrm{d}\rho_{\omega}^{X}(Y),\;\;X\in H_{\mathfrak{c}},\;\;\;\mathbb{P}\mbox{-}\rm{a.s.}$ (67)
@@
    The probability measure $\mu$ therein by the probability measure $\rho$ on $\Omega\times H$ that is naturally associated with the family of random measures $\rho_{\omega}$ and whose marginal on $\Omega$ is given by $\mathbb{P}$; see [117, Prop. 3.6],
@@
    The function space $L^{2}_{\mu}(H_{\mathfrak{c}}\times H_{\mathfrak{f}};H_{\mathfrak{f}})$ therein by the function space $\mathcal{E}=L^{2}_{\rho}(\Omega\times H_{\mathfrak{c}}\times H_{\mathfrak{f}};H_{\mathfrak{f}})$.
By applying to the ambient Hilbert space $\mathcal{E}$, the standard projection theorem onto closed convex sets [123, Theorem 5.2], one defines (given $\Pi_{\mathfrak{c}}$) the conditional expectation $\mathbb{E}_{\rho}[g|\Pi_{\mathfrak{c}}]$ of $g$ as the unique function in $\mathcal{E}$ that satisfies the inequality
@@
$\mathbb{E}_{\rho}[g|\Pi_{\mathfrak{c}}](\omega,X)=\int_{H_{\mathfrak{f}}}g(\omega,X,Y)\,\mathrm{d}\rho_{\omega,X}(Y),$ (71)
@@ IV.2 Non-Markovian optimal reduced model, and conditional expectation
$\mathbb{E}_{\rho_{\theta_{t}\omega}}[G|\Pi_{\mathfrak{c}}](X)=\int_{Y\in H_{\mathfrak{f}}}G(X+Y)\,\mathrm{d}\rho_{\theta_{t}\omega}^{X}(Y),\;X\in H_{\mathfrak{c}},$
@@
$\langle G_{2}(Y,Y)\rangle_{\rho_{\theta_{t}\omega}^{X}}=\int_{Y\in H_{\mathfrak{f}}}\Pi_{\mathfrak{c}}G_{2}(Y,Y)\,\mathrm{d}\rho_{\theta_{t}\omega}^{X}(Y),$ (74)
@@
Hereafter, we simply refer to Eq. (75) as the (random) conditional expectation.
@@ IV.3 Practical implications
Our numerical results demonstrate that this class $\mathcal{P}$ effectively models and approximates the true non-Markovian terms in the optimal reduced model (Eq. (75)) across various scenarios. After learning an optimal parameterization within class $\mathcal{P}$ using a single noise realization, the resulting approximation of the optimal reduced model typically becomes an ODE system with coefficients dependent on the specific noise path. These path-dependent coefficients encode the interactions between noise and nonlinearities, as exemplified by the reduced models Eqns. (85) and (LABEL:Eq_reduced_S_shaped) below, with their ability to predict noise-induced transitions within the reduced state spaces.
@@ V.1 Noise-induced transitions in a stochastic Allen-Cahn model
with the potential $V$ given by the standard double-well function $V(u)=(u^{2}-1)^{2}/4$. Indeed, one can readily check that $-D\mathcal{E}(u)=\partial_{x}^{2}u+u-u^{3}$, where $D\mathcal{E}$ denotes the Fréchet derivative of $\mathcal{E}$ at $u$. The sACE provides a general framework for studying pattern formation and interface dynamics in systems ranging from liquid crystals and ferromagnets to tumor growth and cell membranes [1, 57].
Its universality makes it relevant across diverse fields, contributing to a unified understanding of these complex phenomena. By incorporating the spatio-temporal noise term, $\,\mathrm{d}{\bm{W}_{t}}$, into the deterministic Allen-Cahn equation, the sACE accounts for inherent fluctuation and uncertainty present in real systems [56]. This allows for a more realistic description of transitions between phases, providing valuable insights beyond what pure deterministic models can offer. Near critical points, where phases coexist, the sACE reveals the delicate balance between deterministic driving forces and stochastic fluctuations. Studying these transitions helps us understand critical phenomena in diverse systems, from superfluid transitions to magnetization reversal in magnets [1, 56].
@@
Despite its importance in numerous physical applications, the reduction problem for the accurate reproduction of transition paths in the sACE has received limited attention, with existing works focusing primarily on scenarios near instability onset and low noise intensity (e.g., [132, 103, 102, 133, 134]).
In this study, we consider the sACE placed away from instability onset, after several bifurcations have taken place where multiple steady states coexist; see Fig. 2. It corresponds to the parameter regime
@@
More specifically, the domain size $L$ is chosen so that when $\sigma=0$ the trivial steady state has experienced three successive bifurcations leading thus to a total of seven steady states with two stable nodes ($\phi_{1}^{+}$ and $\phi_{1}^{-}$) and five saddle points (00, $\phi_{2}^{a}$, $\phi_{2}^{b}$, $\phi_{3}^{a}$, and $\phi_{3}^{b}$). These steady states are characterized here by their zeros: the saddle states are of sign-change type while the stable ones are of constant sign; see [126, 135]. Note also that the same conclusion holds for the case with either Neumann boundary conditions or periodic boundary conditions; see e.g. [129, Figure 2]. In our notation, $a$ and $b$ denote the string made of sign changes over $(0,L)$, with leftmost symbol in $a$ to be $+$. For instance, for $\phi_{3}^{a}$ (resp. $\phi_{3}^{b}$), $a=\{+,-,+\}$ (resp. $b=\{-,+,-\}$), while $a=\{+,-\}$ (resp. $b=\{-,+\}$) for $\phi_{2}^{a}$ (resp. $\phi_{2}^{b}$).
@@
Figure 3: Typical and rare transition paths for Eq. (76). One path connecting $u_{0}=0$ to a sign-changing metastable state (typical, left panel), and one path connecting $u_{0}=0$ to a constant-sign metastable state (rare event, right panel). Are shown also the large ($u_{\mathfrak{c}}$) and small ($u_{\mathfrak{f}}$) scale contributions.
Noteworthy is the visual rendering of the sojourn time of the solution path shown in Fig. 2B. It is more substantial in the neighborhood of $\phi_{2}^{a}$ than in that of $\phi_{1}^{+}$. It is a depiction of a solution path to Eq. (76) over $[0,T]$: most of the trajectory spends time near a sign-change metastable state ($\phi_{2}^{a}$) rather than to the constant-sign one ($\phi_{1}^{+}$). Large ensemble simulations of the sACE for the parameter regime (78) reveals that what is shown for a single path here is actually observed at the final time $t=T$, when the noise path is varied: the sign-change states $\phi_{2}^{a/b}$ are the rule whereas the constant-sign states ϕ1±superscriptsubscriptitalic-$\phi_{1}^{\pm}$ are the exceptions, corresponding here to rare events. Figure 3 shows two solution paths in the space-time domain connecting $u_{0}=0$ to either a sign-changing metastable state (typical) or a constant-sign metastable state (rare event). In Section V.3 below we shall discuss the mechanisms behind these patterns and their transition phenomenology, and the extent to which non-Markovian reduced models described below are able to reproduce it.